OpenStudy (anonymous):

help! the measurements of the base and altitude of a triangle are found to be 37 and 32 centimeters. the possible error in each measurement is 0.15. use differentials to estimate the propagated error in computing the area of the triangle. round result to 4 decimal plaaces

OpenStudy (perl):

Area = 1/2 base * height

OpenStudy (anonymous):

oh k

OpenStudy (perl):

A = 1/2*b*h dA = 1/2( b * dh + db * h )

OpenStudy (anonymous):

whats the dh and db suppose to be?

OpenStudy (perl):

differential of height, and differential of base

OpenStudy (perl):

I used the product rule

OpenStudy (perl):

now you can plug in